OpenStudy (anonymous):

1. The area A = πr2 of a circular oil spill changes with the radius. At what rate does the area change with respect to the radius when r = 3ft? 6π ft2/ft 3π ft2/ft 6 ft2/ft 9π ft2/ft 2. The height of a cylinder with a fixed radius of 6 cm is increasing at the rate of 3 cm/min. Find the rate of change of the volume of the cylinder (with respect to time) when the height is 20cm. 36π 108π 360π None of these

OpenStudy (anonymous):

Also these two! 1. A rectangular box has a square base with an edge length of x cm and a height of h cm. The volume of the box is given by V = x2h cm3. Find the rate at which the volume of the box is changing when the edge length of the base is 10 cm, the edge length of the base is increasing at a rate of 3 cm/min, the height of the box is 5 cm, and the height is decreasing at a rate of 1 cm/min. The volume of the box is decreasing at a rate of 200 cm3/min. The volume of the box is increasing at a rate of 400 cm3/min. The volume of the box is decreasing at a rate of 400 cm3/min. The volume of the box is increasing at a rate of 200 cm3/min. 2. A 15-ft ladder rests against a vertical wall. If the top of the ladder slides down the wall at a rate of 0.33 ft/sec, how fast, in ft/sec, is the bottom of the ladder sliding away from the wall, at the instant when the bottom of the ladder is 9 ft from the wall? Answer with 2 decimal places. Type your answer in the space below. If your answer is a number less than 1, place a leading "0" before the decimal point (ex: 0.35).

OpenStudy (anonymous):

so 1 would be 6π ft2/ft??

OpenStudy (anonymous):

6pi

OpenStudy (surry99):

Lets see your work

OpenStudy (anonymous):

\[A=pir^2\] dA/dt=2pir dA/dt=2(pi)(3) 6(pi)

OpenStudy (surry99):

My bad...for this first question you only needed to use the power rule and calculate dA/dr ...the derivative you calculated is dA/dr

OpenStudy (surry99):

on to the cylinder...first you need to look up the volume of a cylinder

OpenStudy (anonymous):

i am so confused.can you show me steps of number 1?

OpenStudy (surry99):

sure...hang on

OpenStudy (surry99):

A = pi*r^2 agree

OpenStudy (surry99):

you there?

OpenStudy (anonymous):

yes sorry

OpenStudy (surry99):

so by the power rule what is da/dr?

OpenStudy (surry99):

dA/dr...(capital A)

OpenStudy (anonymous):

i dont know

OpenStudy (surry99):

are you taking calculus?

OpenStudy (anonymous):

yes

OpenStudy (surry99):

ok. what does the power rule say?

OpenStudy (anonymous):

i dont remebr

OpenStudy (surry99):

ok I need you to look it up and I will help you through these problems

OpenStudy (surry99):

if y = x^n then dy/dx =?

OpenStudy (anonymous):

i understand this problem. i don't understand the cylinder and the rectangle

OpenStudy (surry99):

ok,,,,what is the equation for the volume of cylinder?

OpenStudy (anonymous):

\[V=pir^2h\]

OpenStudy (surry99):

what is dV/dt assuming height is constant?

OpenStudy (anonymous):

\[dV/dt=\pi2r*dr/dt\]

OpenStudy (surry99):

oops what happened to h?

OpenStudy (anonymous):

dh/dt

OpenStudy (surry99):

no...h is constant

OpenStudy (anonymous):

oh so it just stays h?

OpenStudy (surry99):

Yes...because in this particular problem it is constant so you would treat it just like pi...

OpenStudy (anonymous):

ok

OpenStudy (surry99):

remember the if y = constant*x^n then dy/dx = constant*nx^(n-1)

OpenStudy (surry99):

So what is dV/dt...

OpenStudy (anonymous):

what does dr/dt equal?

OpenStudy (anonymous):

r is 6 but what is dr/dt?

OpenStudy (surry99):

fixed radius of 6 cm is increasing at the rate of 3 cm/min.

OpenStudy (surry99):

so dr/dt = ....

OpenStudy (surry99):

is increasing at the rate of 3 cm/min.

OpenStudy (anonymous):

oh i forgot that part so dr/dt=3?

OpenStudy (surry99):

yes...3 cm/mim

OpenStudy (surry99):

so lets see your expression for dV/dt

OpenStudy (anonymous):

so it should equal 720 pi?

OpenStudy (surry99):

I want ot make sure you have dV/dt correct then we substitute

OpenStudy (anonymous):

dV/dt=pi*h*2r*dr/dt

OpenStudy (surry99):

great!!! now substitute and put in your units where applicable

OpenStudy (anonymous):

dV/dt=(20)*(pi)*(2*6)*(3)

OpenStudy (surry99):

I would suggest you put in your units but that is up to you...

OpenStudy (anonymous):

so answer is 720pi?

OpenStudy (surry99):

numerically ..yes it is 720 pi...but there are units with that answer which can help you understand these rate problems

OpenStudy (surry99):

I will be back on later tonight...see what you can do with the next two problems.

OpenStudy (anonymous):

i only need help with the rectangular box. can we do it real quick?

OpenStudy (surry99):

Hint: V=x^2*h but both x and h are changing as a function of time. So how do you calculate dV/dt...you will need to use the product rule of differentiation

OpenStudy (anonymous):

Thank you i need to go to dinner. will try after dinner

OpenStudy (surry99):

your welcome